The table shows a firm’s marginal costs

Output Marginal cost ($)

1 40

2 30

3 20

4 30

5 40

 

If the Average Fixed Cost of producing 5 units is $6, calculate the total cost of producing 5 units.


1
Expert's answer
2021-12-06T17:27:18-0500

"AFC= \frac{TFC}{Output}"


"6 = \frac{TFC}{5}"


TFC = $30

Thus: for 1 unit of output, MC = 40, thus variable cost = 40;total cost =(40 +30) = $70

for 2 units: MC = 30; TC = (70+30) = $100

for 3 units: MC = 20; TC = (20+100) = $120

for 4 units: MC = 30; TC = (30+120) = $150

for 5 units: MC = 40; TC = (40+150) = $190

Thus, the total cost of producing 5 units is $190